OpenStudy (alprincenofl):

How many 3-digit number can be done from the digits 1,2,3,4,5, and 6 which is even and repeating is not allowed.

OpenStudy (cwrw238):

forget all that!! I must be tired

OpenStudy (cwrw238):

the 3 digit numbers must end in 2, 4 or 6 for each of these numbers another 2 numbers are required from the remaining 5

OpenStudy (cwrw238):

so for the numbers ending in 2 there are 5C2 combinations and the same for those ending in 4 and 6

OpenStudy (cwrw238):

so the total number of 3 digit numbers is 3 * 5C2

OpenStudy (cwrw238):

5C2 = (5*4) / (2*1) = 10

OpenStudy (cwrw238):

required total = 30

OpenStudy (cwrw238):

im having a bad day = that's wrong again - the order of the first 2 numbers matters so its not 5C2 but 5P2 which equals 5*4 = 20 so correct answer is 3*20 = 60